# PSF-Med: Measuring and Explaining Paraphrase Sensitivity in Medical Vision Language Models

### v1

@@ Abstract
Medical Vision Language Models (VLMs) can change their answers when clinicians rephrase the same question, which raises deployment risks. We introduce Paraphrase Sensitivity Failure (PSF)-Med, a benchmark of 19,748 chest X-ray questions paired with about 92,000 meaning-preserving paraphrases across MIMIC-CXR and PadChest. Across six medical VLMs, we measure yes/no flips for the same image and find flip rates from 8% to 58%. However, low flip rate does not imply visual grounding: text-only baselines show that some models stay consistent even when the image is removed, suggesting they rely on language priors. To study mechanisms in one model, we apply GemmaScope 2 Sparse Autoencoders (SAEs) to MedGemma 4B and analyze FlipBank, a curated set of 158 flip cases. We identify a sparse feature at layer 17 that correlates with prompt framing and predicts decision margin shifts. In causal patching, removing this feature’s contribution recovers 45% of the yes-minus-no logit margin on average and fully reverses 15% of flips. Acting on this finding, we show that clamping the identified feature at inference reduces flip rates by 31% relative with only a 1.3 percentage-point accuracy cost, while also decreasing text-prior reliance. These results suggest that flip rate alone is not enough; robustness evaluations should test both paraphrase stability and image reliance.
@@ 1 Introduction
Fourth, we demonstrate preliminary mitigations. Clamping Feature 3818 reduces flip rate by 31% relative with 1.3pp accuracy cost. Prompt normalization adds 21% reduction. We release the benchmark and code.
@@ 2.1 Dataset Construction
We draw from two publicly available chest radiograph datasets representing different clinical populations. MIMIC-CXR (Johnson et al., 2019) provides frontal chest radiographs from Beth Israel Deaconess Medical Center in Boston, with questions derived from VQA-RAD (Lau et al., 2018). PadChest (Bustos et al., 2020) provides radiographs from Hospital San Juan in Valencia, Spain, with questions generated from pathology labels. The combined dataset comprises 5,534 images and 19,748 clinical questions.
@@
For each question, we generate 3–5 semantically equivalent paraphrases using GPT-4 (OpenAI, 2023) with instructions to preserve clinical meaning while varying surface form. Transformation strategies include: lexical substitution (synonym replacement), syntactic restructuring (clause reordering), formality shifts (clinical vs. colloquial), and scope variation (quantifiers like “any” or “significant”).
Semantic Filtering.
We apply two filtering steps. First, we compute BioClinicalBERT (Alsentzer et al., 2019) embeddings and retain pairs with cosine similarity >>0.90. Second, we exclude pairs where semantics invert (e.g., “Is there X?” vs. “Is X absent?”). After filtering: $\sim$}92,000 validated paraphrases (mean 4.7 per question).
Table 1: PSF-Med dataset statistics. The benchmark spans two clinical populations with diverse question types and paraphrase transformations.
MIMIC-CXR PadChest Total
Images 1,000 4,534 5,534
Questions 4,998 14,750 19,748
Paraphrases 23,241 68,759 92,000
@@ 2.3 Results: Paraphrase Sensitivity Varies Widely
We evaluate six open-source medical VLMs: MedGemma-4B, MedGemma-27B, and MedGemma-1.5-4B from Google (Google Health AI, 2025); LLaVA-Rad based on BiomedCLIP (Zambrano Chaves and others, 2024); RadFM (Wu et al., 2023); and CheXagent (Chen et al., 2024). All models use greedy decoding with temperature zero to ensure reproducibility.
Table 2: Flip rates (%) across models and datasets. Lower values indicate greater consistency. MedGemma-27B achieves the lowest flip rates; RadFM and LLaVA-Rad show the highest sensitivity.
Model MIMIC-CXR PadChest
MedGemma-27B 8.1 9.9
MedGemma-1.5-4B 9.3 26.8
MedGemma-4B 15.6 42.4
CheXagent 32.6 29.2
LLaVA-Rad 35.9 58.2
RadFM 55.1 58.0
Table 2 presents flip rates across models and datasets. Three main findings emerge from these results.
First, flip rates vary by a factor of seven across models tested on the same benchmark. MedGemma-27B achieves 8.1% on MIMIC-CXR while RadFM reaches 55.1%. This large variation indicates that paraphrase sensitivity is not an inherent property of medical VLMs as a class, but rather depends substantially on specific model architecture, training data, and optimization choices. The implication for practitioners is that model selection matters enormously for deployment in settings where users may phrase questions in varied ways.
@@
Third, robustness does not transfer uniformly across clinical populations. Most models show substantially higher flip rates on PadChest compared to MIMIC-CXR. This likely reflects distribution shift between the US and Spanish healthcare contexts, including differences in imaging protocols, patient demographics, and the language patterns used in clinical annotations. MedGemma-27B is a notable exception, maintaining similar performance across both datasets, which suggests that larger models may generalize more reliably across clinical settings.
@@
Lexical substitutions show the lowest flip rates at just 6–8%. These involve synonym replacement without structural changes, such as “pneumothorax” to “collapsed lung” or “cardiomegaly” to “enlarged heart”. Syntactic restructuring falls in between at 10–15%, covering transformations like active-to-passive voice or clause reordering. This pattern holds consistently across all six models we tested.
@@ 4.3 Semantic Characterization of Feature 3818
Question Phrasing Feature 3818 Response Tendency
@@ 4.5 Toward Mitigation
+ Feature 3818 clamp 10.8% 76.9% 63.1% 32.4%
+ Prompt normalization 12.4% 77.6% 65.9% 30.2%
@@ 6 Conclusion
PSF-Med reveals that paraphrase sensitivity is common in medical VLMs. Flip rates range from 8% to 58% across six models and two datasets. But we also found that low flip rates doesnt mean a model is actually using the image. Text-only baselines show that consistent models often ignore visual input entirely.
@@ A.1 Question Type Distribution
Table 9: Distribution of question types in PSF-Med.
Question Type MIMIC-CXR PadChest Total %
Presence (Is there X?) 2,499 7,375 9,874 50.0
Abnormality (What abnormality?) 1,000 2,950 3,950 20.0
Location (Where is X?) 750 2,213 2,963 15.0
View (Is this frontal?) 499 1,475 1,974 10.0
Severity (Is X mild/severe?) 250 737 987 5.0
Total 4,998 14,750 19,748 100.0
@@ A.2 Paraphrase Type Distribution
Table 10: Distribution of paraphrase transformation types.
@@
Lexical substitution 23,000 25.0
Syntactic restructuring 21,160 23.0
Formality shift 18,400 20.0
Negation-related 16,560 18.0
Scope/quantifier variation 12,880 14.0
Total 92,000 100.0

### v2

@@ Abstract
Medical Vision Language Models (VLMs) can change their answers when clinicians rephrase the same question, a failure mode that threatens deployment safety. We introduce PSF-Med, a benchmark of 26,850 chest X-ray questions paired with 92,856 meaning-preserving paraphrases across MIMIC-CXR, PadChest, and VinDr-CXR, spanning clinical populations in the US, Spain, and Vietnam. Every paraphrase is validated by an LLM judge using a bidirectional clinical entailment rubric, with 91.6% cross-family agreement. Across nine VLMs, including general-purpose models, we find flip rates from 3% to 37%. However, low flip rate does not imply visual grounding: text-only baselines show that some models stay consistent even when the image is removed, suggesting they rely on language priors. To study mechanisms in one model, we apply GemmaScope 2 Sparse Autoencoders (SAEs) to MedGemma 4B and analyze FlipBank, a curated set of 158 flip cases. We identify a sparse feature at layer 17 that correlates with prompt framing and predicts decision margin shifts. In causal patching, removing this feature’s contribution recovers 45% of the yes-minus-no logit margin on average and fully reverses 15% of flips. Acting on this finding, we show that clamping the identified feature at inference reduces flip rates by 31% relative with only a 1.3 percentage-point accuracy cost, while also decreasing text-prior reliance. These results suggest that flip rate alone is not enough; robustness evaluations should test both paraphrase stability and image reliance.
@@ 1 Introduction
Fourth, we demonstrate preliminary mitigations. Clamping Feature 3818 reduces flip rate by 31% relative with 1.3pp accuracy cost. Prompt normalization adds 21% reduction. Concurrent work reinforces why this evaluation axis matters: MIRAGE (Asadi et al., 2026) shows a text-only model can top the ReXVQA benchmark without seeing images, and CheXOne (Zhang et al., 2026) reports high “self-consistency” but measures only stochastic sampling stability, not input rephrasing. We release the benchmark and code at https://huggingface.co/datasets/saillab/psf-med.
@@ 2.1 Dataset Construction
We draw from three publicly available chest radiograph datasets spanning different clinical populations. MIMIC-CXR (Johnson et al., 2019) provides frontal chest radiographs from Beth Israel Deaconess Medical Center in Boston, with questions derived from VQA-RAD (Lau et al., 2018). PadChest (Bustos et al., 2020) provides radiographs from Hospital San Juan in Valencia, Spain. VinDr-CXR (Nguyen et al., 2022) provides radiographs from hospitals in Hanoi, Vietnam. Questions are derived from pathology labels and radiologist annotations. The combined dataset comprises 9,785 images and 26,850 clinical questions.
@@
For each question, we generate 5–6 semantically equivalent paraphrases using GPT-4 (OpenAI, 2023) with instructions to preserve clinical meaning while varying surface form. Transformation strategies include lexical substitution, syntactic restructuring, negation patterns, scope quantification, and specificity modulation.
LLM-Judge Validation.
We validate every question-paraphrase pair using an LLM-as-judge pipeline (Zheng et al., 2023). GPT-5-mini evaluates each pair against a 10-category clinical equivalence rubric testing bidirectional entailment: a paraphrase is kept only if the original and candidate would have the same yes/no answer for every possible chest X-ray. Cross-family validation with Claude Haiku 4.5 shows 91.6% agreement (458/500 pairs). After filtering, we retain 92,856 validated paraphrases (mean 3.5 per question). Negation-pattern paraphrases are rejected at 93.3% and separated into an adversarial reframing slice.
Table 1: PSF-Med v2 dataset statistics after LLM-judge filtering. The benchmark spans three clinical populations with validated paraphrases.
MIMIC PadChest VinDr Total
Images 999 4,523 4,263 9,785
Questions 3,266 14,935 8,649 26,850
Paraphrases 8,938 59,573 24,345 92,856
Para/question 2.7 4.0 2.8 3.5
@@ 2.3 Results: Paraphrase Sensitivity Varies Widely
We evaluate nine VLMs spanning medical and general-purpose models: MedGemma-4B, MedGemma-27B, and MedGemma-1.5-4B from Google (Google Health AI, 2025); two targeted fine-tunes of MedGemma-4B (Targeted LoRA on layers 15–19 and Full LoRA on all 34 layers); LLaVA-Rad (Zambrano Chaves and others, 2024); CheXOne (Zhang et al., 2026); RadFM (Wu et al., 2023); and general-purpose GPT-5-mini and Claude Haiku 4.5. Local models use greedy decoding at temperature zero; API models use deterministic settings.
Table 2: Accuracy (%) and per-paraphrase flip rates (%) on judge-filtered paraphrases. Lower flip rates indicate greater consistency. Results marked use 500 hardest questions; denotes degenerate yes-bias.
MIMIC PadChest VinDr
Model Acc Flip Acc Flip Acc Flip
Targeted LoRA 89.9 4.2 82.1 7.5 74.7 8.5
Full LoRA 84.3 3.3 40.0 10.3 53.5 5.8
MedGemma-4B 88.3 6.7 59.0 13.1 57.2 12.7
MedGemma-27B 76.7 10.6 55.3 17.8 57.6 12.4
MedGemma-1.5-4B 11.5 73.2 12.0 63.8 12.0
CheXOne 65.4 13.3 61.5 10.3 47.3 14.7
LLaVA-Rad 82.9 9.9 83.9 12.4
RadFM 36.8 27.0
GPT-5-mini 52.0 19.3 5.6 8.7 9.2 11.8
Claude Haiku 62.8 22.8
Table 2 presents results across models and datasets. Four findings stand out.
@@
[…]
Third, robustness degrades under distribution shift, but not uniformly. Most models show higher flip rates on PadChest and VinDr compared to MIMIC. Targeted LoRA degrades gracefully (4.2% $\to$ 7.5% $\to$ 8.5%) while MedGemma-27B degrades sharply (10.6% $\to$ 17.8%).
Fourth, general-purpose models are not immune. On the 500 hardest questions (selected by cross-model flip count), GPT-5-mini and Claude Haiku 4.5 show 19.3% and 22.8% flip rates, despite higher baseline accuracy on easier questions in prior evaluations. Even the most capable general-purpose VLMs flip answers when questions are rephrased in clinically equivalent ways.
@@
Lexical substitutions show the lowest flip rates at just 6–8%. These involve synonym replacement without structural changes, such as “pneumothorax” to “collapsed lung” or “cardiomegaly” to “enlarged heart”. Syntactic restructuring falls in between at 10–15%, covering transformations like active-to-passive voice or clause reordering. This pattern holds consistently across all models we tested.
@@ 4.3 Semantic Characterization of Feature 3818
Question Phrasing Feat. 3818 Response
@@ 4.5 Toward Mitigation
+ Feat. 3818 clamp 10.8% 76.9% 63.1% 32.4%
+ Prompt norm. 12.4% 77.6% 65.9% 30.2%
@@ 6 Conclusion
PSF-Med reveals that paraphrase sensitivity is common in medical VLMs. Flip rates range from 3% to 37% across nine models and three datasets. But we also found that low flip rates don’t mean a model is actually using the image. Text-only baselines show that consistent models often ignore visual input entirely.
@@ A.1 Question Type Distribution
Table 9: Distribution of question types in PSF-Med v2 (judge-filtered).
Question Type MIMIC PadChest VinDr Total %
Presence (Is there X?) 956 14,935 5,755 21,646 80.6
Abnormality 850 850 3.2
Location (Where is X?) 345 2,894 3,239 12.1
View (Is this frontal?) 366 366 1.4
Level/Type/Other 749 749 2.8
Total 3,266 14,935 8,649 26,850 100.0
@@ A.2 Paraphrase Type Distribution
Table 10: Distribution of paraphrase transformation types (after LLM-judge filtering).
@@
Lexical substitution 26,850 28.9
Syntactic restructuring 22,048 23.7
Negation pattern 16,714 18.0
Scope/quantifier variation 14,853 16.0
Specificity modulation 12,391 13.3
Total 92,856 100.0